OpenStudy (anonymous):

Get a medal :3 Joby's plant was 6 1/3 inches tall. Pat's was 4 1/2 inches tall. How much taller was Joby's plant than Pat's? A. 1 1/6 in. B. 1 5/6 in. <--- Pretty sure it's this answer C. 2 1/6 in. D. 2 5/6 in. Just want to make sure :3

OpenStudy (anonymous):

6+(1/3) = (19/3) 4+(1/2) = (9/2) (19/3) = (38/6) (9/2) = (27/6) (38/6) - (27/6) = (11/6)

OpenStudy (anonymous):

(11/6) = 1+(5/6)

OpenStudy (anonymous):

So you are correct

OpenStudy (anonymous):

@larryboxaplenty Thanks ^.^